OpenStudy (anonymous):

Please verify, f(x) = x^3-4x^2+4x-16 X^2(x-4) 4(x-4) (X^2+1)(x-4) (x-4) Zeros are: -1,1,4

OpenStudy (the_fizicx99):

You want to find the roots of f(x)? (zero's)

OpenStudy (anonymous):

Yes

OpenStudy (the_fizicx99):

f(x) = x^3 - 4x^2 + 4x - 16 0 = x^3 - 4x^2 + 4x - 16 x^2(x - 4) 4(x - 4) (x-4)(x^2 + 4), x -4 = 0 +4 +4 x = 4

OpenStudy (the_fizicx99):

The sum of two squares (without a GCF) is always prime, http://prntscr.com/39q1av